OpenStudy (anonymous):

solve: sinx (1+cos2x/4)

OpenStudy (ineedhelplz):

Welcome to Open Study! :)

KjSaif (kjsaif):

First Distribute \[\sin{x}+\sin{x}\times \frac{\cos{2x}}{4}\]

KjSaif (kjsaif):

\[\sin{x}\times \frac{\cos{2x}}{4}\] to \[\frac{\sin{x}\cos{2x}}{4}\]

KjSaif (kjsaif):

\[\sin{x}+\frac{\sin{x}\cos{2x}}{4}\]